# Ehrhart quasi-polynomials via Barnes polynomials and discrete moments of parallelepipeds

### v1

@@ Abstract
We also give similar formulas for discrete moments of rational polytopes, and their positive dilates, objects known in the literature as sums of polynomials over a polytope. The appearance of the Barnes polynomials and the Barnes numbers allow for explicit computations. From this work, it is clear that the complexity of computing Ehrhart quasi-polynomials lies mainly in the computation of various discrete moments of parallelepipeds. These discrete moments are in general summed over a particular lattice flow on a closed torus, defined in this paper.
Some of the consequences involve novel vanishing identities for rational polytopes, novel formulations of Ehrhart polynomials of unimodular polytopes, and a differential equation that extends the work of Eva Linke.
@@ 1 Introduction, and statements of the main results
We find here two main objects that form the complete building blocks for the Ehrhart quasi-polynomials of rational polytopes: the collection of Barnes polynomials $B_{k}(t,{\mathbf{a}})$, and the discrete moments of fundamental parallelepipeds, defined in (12) below.
The Barnes polynomials are defined by their generating function:
@@
We use elementary techniques, and our first main result is an explicit formula for the Ehrhart polynomial (resp. quasi-polynomial) of any simple integer (resp. rational) polytope $\mathcal{P}$}, in terms of Barnes polynomials and discrete moments of parallelepipeds. In this sense, this paper also initiates the study of discrete moments of parallelepipeds and more generally the discrete moments of rational polytopes. Such discrete moments may be thought of as belonging to spatial statistics over the integer lattice [13].
For our applications here, we note that if all of the $a_{k}$’s are rational numbers, then definition (1) gives us a polynomial $B_{k}(t,{\mathbf{a}})$ function of $t$, with rational coefficients (see (33) below).
@@
which is a vector whose $k$’th coordinate is the inner product $\langle w_{k},z\rangle$.
The starting point for some of our proofs is Brion’s fundamental result for the integer point transform of a rational polytope $\mathcal{P}$}, defined by
@@
Throughout, we will state our theorems with the words “for almost all $z\in\mathbb{C}^{d}$” to mean that they hold for all $z$ in the set (7) above. We also recall that a quasi-polynomial is a function $f(t)=\sum_{k=0}^{d}c_{k}(t)t^{k}$, where each coefficient $c_{k}(t)$ is a periodic function of $t\in\mathbb{Z}_{>0}$. Throughout, even after we extend the domain of $c_{k}(t)$ and hence $f(t)$ to all positive real dilations $t\in\mathbb{R}_{>0}$, as in Linke’s paper [28], we will continue to call $f(t)$ a quasi-polynomial.
@@
where the $B_{j}$’s are the classical Bernoulli numbers [7, Theorem 2]. In section 2 below, we give some additional background and examples of Barnes polynomials and Barnes numbers.
The Barnes numbers extend the usual Bernoulli numbers, in the precise sense that for dimension $d=1$, and for ${\mathbf{a}}=(1)$, we have $B_{j}({\mathbf{a}})=B_{j}$, the $j$’th Bernoulli number. A very useful explicit polynomial representation for the Barnes polynomials, in terms of the Barnes numbers $B_{k}({\mathbf{a}})$ is:
@@
as in Figure 2, Section 4 below. We will compute in a straightforward way some explicit lattice flows over half-open parallelepipeds (see Sections 3 and 4).
@@
for each fixed integer $k\geq 0$, $v\in\mathbb{R}^{d}$ and $t\geq 0$.
@@
It may seem rather surprising a priori that the right-hand-sides of both (16) and (17) appears to depend on $z\in\mathbb{C}^{d}$; but in fact a posteriori they are independent of $z$, as the left-hand-side of (16) shows. Moreover, it is possible that judicious choices of $z$ may reduce the computational complexity of the ensuing discrete moments, and therefore of the Ehrhart polynomials and quasi-polynomials.
@@
for all $t>0$. We remark that prima facie these discrete moments may not appear to be polynomial-time computable. However, see Conjecture 1 below.
@@
We quickly recover a couple of the main results of Eva Linke’s seminal paper [28]. Linke’s important paper opened the door to the study of all positive real dilates of rational polytopes. Theorem 1.5 of [28] proves that the coefficients of quasi-polynomials of rational polytopes are always piecewise polynomial functions of $t$. Here it follows easily from Theorem 1. We may rewrite (15) as follows:
@@
recovering one of the other main results in Linke’s paper, namely [28, Theorem 1.6].
@@
When the vertices are rational points $v\in\mathbb{Q}^{d}$, and we let the lattice flow proceed for all sufficiently large positive values of $t$, we topologically get a finite collection of closed geodesics on the torus, where each geodesic is an orbit of one integer point in $\mathcal{P}$}. As we see from Theorem 1, the periodicity of the Ehrhart quasi-polynomials arises precisely because of the periodicty of these closed geodesics.
@@
The ODE (20) in Theorem 2, part (c), directly extends Linke’s ODE in her seminal paper [28]. It is amusing - and a bit surprising - that we obtain exactly the same ODE, even though our ODE describes discrete moments of dilated rational polytopes.
@@
Let ${\mathcal{P}}\subset\mathbb{R}^{d}$ be a simple $d$-dimensional integer polytope. Then $\mathcal{P}$} possesses the following set of finite vanishing identities, valid for almost all $z\in\mathbb{C}^{d}$:
@@
    $0$ $=\sum_{k=0}^{m}\binom{m}{k}\sum_{v\in V}B_{k}\Big(t\langle v,z\rangle,{\mathbf{a}}_{v}\Big)\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}{\langle p,z\rangle}^{m-k}.$ (21)
@@
    $0$ $=\sum_{j=0}^{m-r}\binom{m-r}{j}\sum_{v\in\rm{V}}\langle v,z\rangle^{r}B_{j}({\mathbf{a}}_{v})\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{d}}{\langle p,z\rangle}^{m-r-j}.$ (22)
@@
For example, a special case of Theorem 3, for $m=1,r=0$, as we show below in (106), is:
@@
allowing us to rapidly compute certain weighted linear combinations of discrete moments of fundamental parallelepipeds, over all vertices of $\mathcal{P}$}. The cases with $r=m$ appeared in [12], while all the other cases of Theorem 3appear to be new. We expand on these special cases in (101) below. See also [37, Theorem 5.11] for a Fourier-based proof of these special cases.
A unimodular polytope is a simple, integer $d$-dimensional polytope ${\mathcal{P}}\subset\mathbb{R}^{d}$ such that at each vertex $v\in{\mathcal{P}}$, there is a collection of edge vectors $w_{1}(v),\dots w_{d}(v)$ that forms a basis for $\mathbb{Z}^{d}$. Unimodular polytopes are sometimes called “smooth polytopes” in toric geometry, and “delzant polytopes” in symplectic geometry. It is well known that in practice, there are numerous results - regarding the discrete geometry of unimodular polytopes - which are easy to prove. From the present perspective, this ease is due to the fact that most of the discrete moments at their vertices vanish, as we observe next, leading to an extreme simplification of the main results.
Corollary 2.
Let ${\mathcal{P}}\subset\mathbb{R}^{d}$ be a $d$-dimensional unimodular polytope. Then we have:
$L_{\mathcal{P}}(t)=\frac{(-1)^{d}}{d!}\sum_{j=0}^{d}\left(\binom{d}{j}\sum_{v\in V}B_{j}({\mathbf{a}})\langle v,z\rangle^{d-j}\right)t^{d-j},$ (24)
(Proof) $\square$
Now we consider ${\mathcal{P}}^{o}$, the interior of the polytope $\mathcal{P}$}. Ehrhart’s reciprocity law, namely that $L_{{\mathcal{P}}^{o}}(t)=(-1)^{d}L_{{\mathcal{P}}}(-t)$, gives us the following immediate corollary of Theorem 1 for open, rational simple polytopes. All of the results above easily carry over to open polytopes, with their open vertex tangent cones. The only difference is the consideration of the interior of a fundamental parallelepiped, which is defined by:
$\text{int}(\Pi_{v}):=\{\sum_{k=1}^{d}\lambda_{k}w_{k}(v)\mid 0<\lambda<1\},$ (25)
for each vertex $v\in{\mathcal{P}}$. In all of the corollaries, we keep using the same notation as above for $B_{d}(x,{\mathbf{a}})$, $B_{j}({\mathbf{a}})$, and ${\mathbf{a}}_{v}$.
Corollary 3.
Let ${\mathcal{P}}\subset\mathbb{R}^{d}$ be a $d$-dimensional simple rational polytope. Let ${\mathcal{P}}^{o}$ be the corresponding open polytope, and let $\text{int}(\Pi)$ be the open parallelepiped at the vertex $v\in{\mathcal{P}}$.
  1. (a)
    The quasi-polynomial of ${\mathcal{P}}^{o}$ is equal to:
    $L_{{\mathcal{P}}^{o}}(t)=\frac{1}{d!}\sum_{v\in V}\sum_{k=0}^{d}\binom{d}{k}B_{k}\big(-t\langle v,z\rangle,{\mathbf{a}}_{v}\big)\sum_{p\in\mathbb{Z}^{d}+tv\;(\hskip-2.84544pt\bmod\text{int}(\Pi_{v}))}{\langle p,z\rangle}^{d-k},$ (26)
    valid for almost all $z\in\mathbb{C}^{d}$, and all $t>0$.
  2. (b)
    $b_{r}(t)=\frac{(-1)^{r}}{d!}\sum_{v\in V}\langle v,z\rangle^{r}\sum_{j=0}^{d-r}\binom{d}{r,j}B_{j}({\mathbf{a}}_{v})\sum_{p\in\mathbb{Z}^{d}+tv\;(\hskip-2.84544pt\bmod\text{int}(\Pi_{v}))}{\langle p,z\rangle}^{d-r-j},$ (27)
    valid for almost all $z\in\mathbb{C}^{d}$, and all $t>0$. $\square$
Remark 6.
A lattice polytope is called hollow if its interior contains no integer points. There is considerable current research on classifying hollow simplices [1]. An immediate consequence of the latter corollary is the following equivalence for this property.
Corollary 4.
for almost all $z\in\mathbb{C}^{d}$. $\square$
We also get the following interesting identities that arise from the standard topological fact that the constant term of the Ehrhart polynomial of an integer polytope is always $1$.
@@
Specializng Corollary 5 to unimodular polytopes, we have:
$1=\frac{(-1)^{d}}{d!}\sum_{v\in V}B_{k}({\mathbf{a}}_{v}),$ (31)
and the proof is identical to the proof of Corollary 2. For example, it follows from (31) that for integer unimodular polygons, we have the amusing identity:

### v2

@@ Abstract
We also give similar formulas for discrete moments of rational polytopes, and their positive dilates, objects known in the literature as sums of polynomials over a polytope. The appearance of the Barnes polynomials and the Barnes numbers allow for explicit computations. From this work, it is clear that the complexity of computing Ehrhart quasi-polynomials lies mainly in the computation of various discrete moments of parallelepipeds. These discrete moments are in general summed over a particular lattice flow on a compact torus, defined in this paper.
Some of the consequences involve novel vanishing identities for rational polytopes. As another consequence, we obtain a differential equation for discrete moments of rational polytopes, which extends the work of Eva Linke. For smooth polytopes, we obtain novel and much simpler formulations of Ehrhart polynomials, discrete moments, and vanishing identities that may be of independent interest from the perspective of Barnes polynomials and Barnes numbers. These formulations show the utility of Barnes polynomials in geometric combinatorics, due to their very rich structure that extends the $1$-dimensional Bernoulli polynomials.
@@ 1 Introduction, and statements of the main results
We find here two main objects that form the complete building blocks for the Ehrhart quasi-polynomials of rational polytopes: the collection of Barnes polynomials $B_{k}(t,{\mathbf{a}})$, and the discrete moments of fundamental parallelepipeds, defined in (12) below. The Barnes polynomials are defined by their generating function:
@@
We use elementary techniques, and our first main result is an explicit formula for the Ehrhart quasi-polynomial of any simple integer rational polytope $\mathcal{P}$}, in terms of Barnes polynomials and discrete moments of parallelepipeds. In this sense, this paper also initiates the study of discrete moments of half-open parallelepipeds and more generally the discrete moments of rational polytopes. Such discrete moments may be thought of as belonging to spatial statistics over the integer lattice [14].
@@
which is a vector whose $k$’th coordinate is the inner product $\langle w_{k},z\rangle$. The starting point for some of our proofs is Brion’s fundamental result for the integer point transform of a rational polytope $\mathcal{P}$}, defined by
@@
Throughout, we state our theorems with the words “for almost all $z\in\mathbb{C}^{d}$” to mean that they hold for all $z$ in the set (7) above. We also recall that a quasi-polynomial is a function $f(t)=\sum_{k=0}^{d}c_{k}(t)t^{k}$, where each coefficient $c_{k}(t)$ is a periodic function of $t\in\mathbb{Z}_{>0}$. Throughout, even after we extend the domain of $c_{k}(t)$ and hence $f(t)$ to all positive real dilations $t\in\mathbb{R}_{>0}$, as in Linke’s paper [32], we will continue to call $f(t)$ a quasi-polynomial.
@@
where the $B_{j}$’s are the classical Bernoulli numbers [8, Theorem 2]. Equation (9) also shows that $(a_{1}\cdots a_{d})B_{k}({\mathbf{a}})$ is a homogeneous polynomial of the $a_{j}$’s, of degree $k$. Appendix A gives some additional background and examples of Barnes polynomials and Barnes numbers. The Barnes numbers extend the usual Bernoulli numbers, in the precise sense that for dimension $d=1$, and for ${\mathbf{a}}=(1)$, we have $B_{j}({\mathbf{a}})=B_{j}$, the $j$’th Bernoulli number. In many of our results we will use the following explicit polynomial representation for the Barnes polynomials in terms of the Barnes numbers $B_{k}({\mathbf{a}})$:
@@
as in §2.2, Figure 2 below. We will compute in a straightforward way some explicit lattice flows over half-open parallelepipeds (see §2.1 and §2.2).
@@
for each fixed integer $k\geq 0$, $v\in\mathbb{R}^{d}$ and $t\in\mathbb{R}$.
@@
It may seem rather surprising a priori that the right-hand-sides of both (13) and (14) appears to depend on $z\in\mathbb{C}^{d}$; but in fact a posteriori they are independent of $z$, as the left-hand-side of (13) shows. Moreover, it is possible that judicious choices of $z$ may reduce the computational complexity of the ensuing discrete moments, and therefore of the Ehrhart polynomials and quasi-polynomials (see §8, Remark 10).
@@
for all $t>0$. We remark that prima facie these discrete moments may not appear to be polynomial-time computable. However, Barvinok’s algorithm [6] shows that they are (see also Question 1 below).
@@
We quickly recover a couple of the main results of Eva Linke’s important paper [32]. Linke’s work opened the door to a more precise study of all positive real dilates of rational polytopes. Theorem 1.5 of [32] proves that the coefficients of quasi-polynomials of rational polytopes are always piecewise polynomial functions of $t$. Here it follows easily from Theorem 1 by rewriting (15) as follows:
@@
recovering one of the other main results in Linke’s paper, namely [32, Theorem 1.6].
@@
Given a rational vertex $v\in\mathbb{Q}^{d}$, let’s allow the lattice flow $\mathbb{Z}^{d}-tv\pmod{\Pi}$ to proceed for all sufficiently large positive values of $t$. Topologically, we get a finite collection of closed geodesics on the torus, where each geodesic is an orbit of one integer point in $\mathcal{P}$}. As we see from Theorem 1, the periodicity of the Ehrhart quasi-polynomials arises precisely because of the periodicity of these closed geodesics, appearing in the discrete moments of the fundamental parallelepipeds of the vertices.
@@
Remark 5.
We observe that Theorem 1 also implies that for an integer polytope $\mathcal{P}$}, allowing all positive dilates $t>0$ quickly implies here that the periodic components always have a period of $1$, because for integral vertices $v\in{\mathcal{P}}$, the domain of summation in the discrete moments of formula (13) satisfies $\mathbb{Z}^{d}-v=\mathbb{Z}^{d}$. It becomes an interesting question of whether there exists a smaller period, $0<t<1$, for some integer polytopes.
@@
The ODE (20) in Theorem 2, part (c), directly extends Linke’s ODE in her seminal paper [32]. It is amusing - and a bit surprising - that we obtain exactly the same ODE as in Linke’s work, even though our ODE describes discrete moments of dilated rational polytopes.
1.1 Vanishing identities for rational polytopes
@@
Let ${\mathcal{P}}\subset\mathbb{R}^{d}$ be a simple $d$-dimensional rational polytope. Then $\mathcal{P}$} possesses the following set of finite vanishing identities, valid for almost all $z\in\mathbb{C}^{d}$ and for all $t>0$:
@@
    $0$ $=\sum_{k=0}^{m}\binom{m}{k}\sum_{v\in V}B_{k}\Big(t\langle v,z\rangle,{\mathbf{a}}_{v}\Big)\sum_{q\in\mathbb{Z}^{d}-tv\;(\hskip-2.84544pt\bmod\Pi_{v})}{\langle q,z\rangle}^{m-k}.$ (21)
@@
    $0$ $=\sum_{v\in\rm{V}}\langle v,z\rangle^{r}\sum_{j=0}^{m-r}\frac{B_{j}({\mathbf{a}}_{v})}{j!(m-r-j)!}\sum_{q\in\mathbb{Z}^{d}-tv\;(\hskip-2.84544pt\bmod\Pi_{v})}{\langle q,z_{0}\rangle}^{m-r-j}.$ (22)
@@
For example, in the special case of integer polytopes, Theorem 3 with $m=1,r=0$ tells us that
@@
as we show below in (94). This allows us to rapidly compute certain weighted linear combinations of discrete moments of fundamental parallelepipeds, over all vertices of $\mathcal{P}$}. The cases with $r=m$ appeared in [13], while all the other cases of Theorem 3 appear to be new. We expand on these special cases $r=m$ in (89) below (See also [42, Theorem 5.11] for a Fourier-based proof).
1.2 The constant term 1 identity
The following interesting identities arise from the standard topological fact that the constant term of the Ehrhart polynomial of an integer polytope is always $1$.
@@ 1.2 The constant term 1 identity
1.3 Open polytopes
Now we consider ${\mathcal{P}}^{o}$, the interior of the polytope $\mathcal{P}$}. Ehrhart’s reciprocity law, namely that $L_{{\mathcal{P}}^{o}}(t)=(-1)^{d}L_{{\mathcal{P}}}(-t)$, gives us the following immediate corollary of Theorem 1 for open, rational simple polytopes. All of the results above easily carry over to open polytopes, with their open vertex tangent cones.
In all of the corollaries, we keep using the same notation as above for $B_{d}(x,{\mathbf{a}})$, $B_{j}({\mathbf{a}})$, and ${\mathbf{a}}_{v}$.
Corollary 3.
Let ${\mathcal{P}}\subset\mathbb{R}^{d}$ be a $d$-dimensional simple rational polytope. Let ${\mathcal{P}}^{o}$ be the corresponding open polytope.
  1. (a)
    The quasi-polynomial of ${\mathcal{P}}^{o}$ is equal to:
    $L_{{\mathcal{P}}^{o}}(t)=\frac{1}{d!}\sum_{v\in V}\sum_{k=0}^{d}\binom{d}{k}B_{k}\big(-t\langle v,z\rangle,{\mathbf{a}}_{v}\big)\sum_{p\in\mathbb{Z}^{d}+tv\;(\hskip-2.84544pt\bmod\Pi_{v})}{\langle p,z\rangle}^{d-k},$ (26)
    valid for almost all $z\in\mathbb{C}^{d}$, and all $t>0$.
  2. (b)
    $b_{r}(t)=\frac{(-1)^{r}}{d!}\sum_{v\in V}\langle v,z\rangle^{r}\sum_{j=0}^{d-r}\binom{d}{r,j}B_{j}({\mathbf{a}}_{v})\sum_{p\in\mathbb{Z}^{d}+tv\;(\hskip-2.84544pt\bmod\Pi_{v})}{\langle p,z\rangle}^{d-r-j},$ (27)
    valid for almost all $z\in\mathbb{C}^{d}$, and all $t>0$. $\square$
@@ 1.3 Open polytopes
A lattice polytope is called hollow if its interior contains no integer points. There is considerable current research on classifying hollow simplices [1]. An immediate consequence of the Corollary 3 is the following equivalence for this property.
Corollary 4.
for almost all $z\in\mathbb{C}^{d}$. $\square$
1.4 Smooth polytopes
A smooth polytope is a simple, integer $d$-dimensional polytope ${\mathcal{P}}\subset\mathbb{R}^{d}$ such that at each vertex $v\in{\mathcal{P}}$, there is a collection of edge vectors $w_{1}(v),\dots w_{d}(v)$ that forms a basis for $\mathbb{Z}^{d}$. The terminology of Smooth polytopes comes from toric geometry, and smooth polytopes are sometimes also called “totally unimodular polytopes” in integer programming, or “delzant polytopes” in symplectic geometry. It is well known that in practice, there are numerous results - regarding the discrete geometry of smooth polytopes - which are easy to prove. From the present perspective, this ease is due to the fact that most of the discrete moments at their vertices vanish, as we observe next, leading to an extreme simplification of the main results.
First, specializing Theorem 1 to smooth polytopes immediately gives us the following formula for their Ehrhart polynomials. The proofs of the rest of the corollaries are omitted, because their immediate proofs are identical to that of Corollary 5.
Corollary 5.
Let ${\mathcal{P}}\subset\mathbb{R}^{d}$ be a $d$-dimensional smooth polytope. Then we have:
$L_{\mathcal{P}}(t)=\frac{(-1)^{d}}{d!}\sum_{j=0}^{d}\left(\binom{d}{j}\sum_{v\in V}B_{j}({\mathbf{a}}_{v})\langle v,z\rangle^{d-j}\right)t^{d-j},$ (29)
for all $t\in\mathbb{Z}_{>0}$, and almost all $z\in\mathbb{C}^{d}$. (Proof) $\square$
Specializing Theorem 2 to smooth polytopes gives the following moment formulas for smooth polytopes.
Corollary 6.
Let ${\mathcal{P}}\subset\mathbb{R}^{d}$ be a $d$-dimensional smooth polytope. For each positive integer $m$, we have the following formula for the $m$’th discrete moment of positive dilates of $\mathcal{P}$}:
$\sum_{p\in t{\mathcal{P}}\cap\mathbb{Z}^{d}}\langle p,z\rangle^{m}$ $=(-1)^{d}\frac{m!}{(d+m)!}\sum_{v\in V}B_{d+m}(t\langle v,z\rangle,{\mathbf{a}}_{v})$ (30)
valid for almost all $z\in\mathbb{C}^{d}$, and all $t\in\mathbb{Z}_{>0}$. $\square$
Specializing Theorem 3 to smooth polytopes immediately gives the following set of vanishing identities.
Corollary 7.
Let ${\mathcal{P}}\subset\mathbb{R}^{d}$ be a $d$-dimensional smooth polytope. Then $\mathcal{P}$} possesses the following set of finite vanishing identities, valid for almost all $z\in\mathbb{C}^{d}$, and for all $t\in\mathbb{Z}_{>0}$:
  1. (a)
    For each $0\leq m\leq d-1$, we have:
    $0$ $=\sum_{v\in V}B_{m}\Big(t\langle v,z\rangle,{\mathbf{a}}_{v}\Big).$ (31)
  2. (b)
    Equivalently, we may also express the latter set of identities by using the Barnes numbers, and removing the variable $t$. For each $(m,r)$ with $0\leq m\leq d-1,\text{ and }0\leq r\leq m$, and for almost all $z\in\mathbb{C}^{d}$ we have:
    $0$ $=\sum_{v\in\rm{V}}\langle v,z\rangle^{r}B_{m-r}({\mathbf{a}}_{v}).$ (32)
    $\square$
Specializing Corollary 2 to smooth polytopes, we have the following.
Corollary 8.
Let ${\mathcal{P}}\subset\mathbb{R}^{d}$ be a $d$-dimensional smooth polytope. The we have:
$1=\frac{(1)^{d}}{d!}\sum_{v\in V}B_{d}({\mathbf{a}}_{v}),$ (33)
for almost all $z\in\mathbb{C}^{d}$. $\square$
For example, it follows from Corollary 8 that for integer smooth polygons, we have the amusing identity: